QUESTION: Calculate the total mass of carbon and hydrogen atoms from Question A and compare that value to the original sample mass given. Upon comparison, do your values make sense? Name the scientific law that supports your reasoning.

Question A said "a 10.00g unknown sample with a formula of CxHY was analyzed by combustion analysis. The sample produced 27.42g of CO2(g) and 22.46g H2O(g). Find the empirical formula of the unknown sample." --> I determined that the empirical formula was CH4
Answer #1

mass of CO2 = 27.42 g

moles of CO2 = 27.42 / 44.01 = 0.623 mol

moles of C = 0.623

mass of C = 0.623 x 12 = 7.4765 g

moles of H2O = 22.46 / 18.02 = 1.246 mol

moles of H = 2 x 1.246 = 2.49

mass of H = 2.49 g

total mass of carbon and hydrogen = 7.4765 + 2.493

total mass of carbon and hydrogen = 9.97 g

mass of sample = 10.00 g

approximately mass of sample and total mass of H and C are equal.

yes , the values make sense